The most recent financial statements for Alexander Co. are shown here:

Income Statement Balance Sheet
  Sales $ 44,000 Current assets $ 18,140 Long-term debt $ 37,420
  Costs 35,900 Fixed assets 68,700 Equity 49,420
  Taxable income $ 8,100   Total $ 86,840   Total $ 86,840
  Taxes (23%) 1,863      
    Net income $ 6,237

Assets and costs are proportional to sales. The company maintains a constant 30 percent dividend payout ratio and a constant debt-equity ratio.

What is the maximum dollar increase in sales that can be sustained assuming no new equity is issued? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Dividend payout is 30%, as Earnings = Retention ratio + dividend ratio,

Retention ratio(b) = 1 - 0.3 = 70%

As growth = b * r, we have to caluclate r, which is ROE,

ROE = Net income / Equity = 6,237 / 49,420 = 12.62%

The sustained increase in sales is sustained in propotion to the increase in growth,

Sustained increase in growth = r * b / 1-(r * b)

=0.1262 * 0.70 / 1- (0.1262 * 0.70)

= 9.690350%

Maximum dollar increase in sales = sales * Sustained increase in growth

=44000 * 9.690350%

=$4263.75